OpenStudy (anonymous):

can some one help me with the question y= x^2+5x+6

OpenStudy (anonymous):

3*2

OpenStudy (anonymous):

okay, but what exactly is the question asking you to do

OpenStudy (anonymous):

graph each parabola and give its vertex , axis, x-intercepts , and y-intercept

OpenStudy (anonymous):

okay, once we factored it we can set y to be 0 and solve for x, thus obtaining the x-intercepts: (x+3)(x+2)=0 x+3=0--->x=-3 x+2=0----->x=-2 x-intercepts at (-3,0) and (-2,0)

OpenStudy (anonymous):

the x-coordinate of the vertex is giveb by -b/2a here b=5 and a=1, thus: -(5)/(2*1)=-5/2, now if we plug this in to our original equation we get: (-5/2)^2+5(-5/2)+6=-1/4, thus the vertex is located at (-5/2,-1/4)

OpenStudy (anonymous):

the y intercept is given by setting x to be zero and solving for y, this : y=(0)^2+5(0)+5 y=6 thus the y intercept is at (0,6)

OpenStudy (anonymous):

thank you and alsoo i have another ques

OpenStudy (anonymous):

f(x)=2x^2+8x-8